OpenStudy (anonymous):

HELP!! A truck rental company rents a moving truck for one day by charging $37 plus $0.10 per mile. Write a linear equation that relates the cost C, in dollars, of renting the truck to the number x of miles driven. what is the cost of renting the truck if the truck is driven 151 miles? 457 miles? type the linear equation that relates the cost C, in dollars of renting the truck to the number of x miles driven. C=_________ what is the cost of renting the truck if the truck is driven 151 miles. C=$_________ What is the cost of renting the truck if the truck is driven 457 miles? c=$___________

OpenStudy (stamp):

k

OpenStudy (stamp):

A truck rental company rents a moving truck for one day by charging $37 plus $0.10 per mile. charging $37 plus $0.10 per mile. c = $37 plus $0.10 per mile. c = $37 + $0.10 per mile. c = $37 + $0.10*mile. c = 37 + 0.10m

OpenStudy (stamp):

cost of renting the truck if the truck is driven 151 miles? c if the truck is driven 151 miles? c if driven 151 miles? c if 151 miles? c if 151m? c = 37 + 0.10m m = 151 c = 37 + 0.10(151) c = 37 + (15.1) c = 52.1

OpenStudy (stamp):

I am not doing the third one for you

OpenStudy (anonymous):

its cool i just realized that i could had done this myself after I reread it again. but thanks!!